Ques:1
1. Find r ′ and T for r = ht 2 , 1, ti
2. Find r ′ and T for r = hcos t, sin 2t, t2 i.
1
Expert's answer
2021-07-09T09:43:04-0400

Solution;

1.Find r' and T for r=ht2,1,ti

r'="\frac {d}{dt}" r=2ht,0,i

T="\frac{r'}{|r'|}"

"|r'|=\sqrt{(2ht)^2+0^2+i^2}=\sqrt{4h^2t^2-1}"

T="\frac{2ht}{\sqrt{4h^2t^2-1}},0,\frac{i}{\sqrt{4h^2t^2-1}}"

2.Find r' and T for r=hcost ,sin2t ,t2i

r'="\frac{d}{dt}"r=-hsint,2cos2t,2it

T="\frac{r'}{|r'|}"

"|r'|=\sqrt{(-hsint)^2+(2cos2t)^2+(2it)^2}" ="\sqrt{h^2sin^2t+4cos^2(2t)-4t^2}"

T="\frac{-hsint}{\sqrt{h^2sin^2t+4cos^2(2t)-4t^2}}," "\frac{2cos2t}{\sqrt{h^2sin^2t+4cos^2(2t)-4t^2}}," "\frac{2it}{\sqrt{h^2sin^2t+4cos^2(2t)-4t^2}}"